# Bubble dynamics in a cavitating venturi

### v1

@@ 2.5.1 Case-1
The void fraction plot, as shown in figure 14, increases continuously up to 3.8 ×$\times 10^{-4}$ after the nozzle exits from the location of 2.7 to the far downstream of the nozzle to reach the value of 1 which corresponds to fully vapor condition. From the plot of Coefficient of Pressure given in the figure 15, the same behaviour is inferred.
@@ 2.5.2 Case-2
Figure 19: Non-dimensional bubble radius plot for the case2 of Liquid Nitrogen
@@ 2.5.3 Case-3
It is concluded that the 1D numerical model for cavitating flows with the thermal effects resulted in successful benchmarking of the Wang &\&& Brennen [9] solution for water. The extension of the same model for cryogenic cavitating nozzle flows resulted in the damping nature of solutions as the behavior of cryogenic liquid nitrogen distinctly differs from the water. Overall, it can be ascertained that the 1D numerical model gives a limited predictability of the nature of cavitating flows, which are restricted to Bubble Radius, Void fractions, Coefficient of Pressure, and the thermal and pressure source terms corresponding to a single bubble as the governing Rayleigh Plesset equation itself is valid for dynamics of the single bubble only. However, cavitation is a cluster phenomenon that demands the extension of the studies to two-dimensional numerical and experimental studies, whose attempted results are given in the subsequent sections.
@@ 3 Two-dimensional model CFD Studies
The following section systematically starts with an overview of the methodology adopted for 2D numerical studies using the commercial CFD solver ANSYS Fluent. The simulations have been carried out to capture the cavitating flows in a Venturi profile chosen to be symmetric, with a turbulence model to get insight into the actual experimental work. Simulations were done for cryogenic liquid nitrogen. The Governing Equations, such as continuity and momentum for the multiphase model, are solved using finite volume fluent code. The current chapter discusses the viscous modeling, interfacial effects, solution methodology, and the boundary conditions used for the benchmarking test case of Hord et al. [5] and the extension of the same for current 2D numerical work.
@@ 3.1 Turbulence Modelling
The realistic cases of cavitating nozzle and venturi flows are turbulent in nature. Hence, a standard turbulence model has been incorporated into the simulations. Realizable $k-\varepsilon$ was used in simulations. The chosen model of $k-\varepsilon$ is in line with the internal flows, and the coefficients for the same are the standard values built into the ANSYS-fluent. The model has two transport equations, one for turbulent kinetic energy k and another for turbulent dissipation $\varepsilon$. The two transport equations are given as follows
@@
The turbulent viscosity $\mu_{t}=\rho C_{\mu}\frac{k^{2}}{\varepsilon}$, where $C_{\mu}$ is a function of mean strain rate and mean rotation rate, the angular velocity of the system rotation, and the turbulence fields such as k and $\varepsilon$. The model constants are given by,
@@ 3.2.1 Schnerr Sauer model for Mass Transfer
Rb is the bubble radius, and $\eta_{b}$ is the bubble population number per unit volume of the liquid. In this model assumes that the number of bubbles is neither created nor destroyed. To model bubble dynamics, the Rayleigh-Plesset equation in simple form was incorporated that was built with Schnerr-Sauer model, which is given by
@@
Where Pb is the bubble pressure, P is the pressure which is felt far away from the bubble. Using the above relations, the mass source term is
@@ 3.3 Interfacial Heat Transfer Effects
3.4 Benchmark of 2D simulations with experimental work of Hord.et al
An attempt was made to benchmark the 2D numerical work involved in this work with the experimental test case of Hord 121B [5] for validating the cavitating length for liquid hydrogen flows in a venturi profile.
A finite volume mesh was done using Ansys for the actual profile used by Hord for his experiment. A grid independence study was done for the Venturi Profile, and the grid independent solution was obtained for a mesh with 51,335 nodes and 50527 elements. The following figure shows the meshed model of the venturi profile.
@@
The simulation was done with the boundary conditions, which were taken from the experimental data of Hord et al. [5].
@@ 3.4.1 Solution Methodology
The simulations were performed using fluent with the Eulerian Multiphase model. An implicit finite volume scheme is based on a segregated pressure-based solver with SIMPLE as the pressure-velocity coupling. A second-order upwind scheme was used for continuity, momentum, energy, turbulent dissipation, and Kinetic energy.
@@
The residuals were set to be 0.001 for continuity, velocity, K, and $\varepsilon$. For Energy and volume fraction, the residuals were set to be 1-06.
@@ 3.4.2 Simulation Results
The following figure32 shows the simulation results of Hord test case 121B. The blue fill indicates the complete liquid phase, and the red fill indicates the complete vapor phase.
@@ 3.4.3 Estimating Cavitating length
Similar work was done by Rodio et al. [8], and the contour of the void fraction of the vapor obtained for the Hord Test case 121B [5] is shown in the following figure.
@@
It was observed from the results that the simulation performed by Rodio et al. [8] overpredicted the experimental cavitating length observed by Hord et al. [5]. The following measurement proofs validate the above interpretation of the result.
@@
The calculated cavitating length from the above contours of volume fraction gives a length of 11.8235 cm against the actual Experimental cavitating length of 8.26 cm, measured by Hord et al. [5]. However, the wrong prediction of cavitation length may not be precisely due to the model used, as mesh methods and wall functions have a great influence on solution convergence. Having benchmarked the 2D numerical model for cryogenic cavitating venturi with Hord et al. [5] experimental test cases for predicting the cavitating length in the venturi profile. The same is now extended for the current work, the details of which are given in the following section.
@@ 4.1 Computational Model Details
A steady state, two-dimensional, incompressible, turbulent Eulerian mixture model was developed based on finite volume formulation. The details of the computational model are shown below
@@
The figure41depicts the actual 2D model of the venturi region of the test section developed for simulations. The dimensions correspond to the actual dimensions of the experimental setup. The boundaries include an inlet, wall, symmetry, and an outlet.
@@ 4.2 Boundary conditions
The outlet boundary condition was atmospheric pressure, as the experiment’s venturi profile opens to the atmosphere. Liquid temperature condition at the outlet is assumed to be 77 K, and vapour fraction at the outlet is assumed to be 0 as the single-phase liquid was desired.
@@ 4.3 Solution Methodology
The same solution methodology was used for the Hord benchmark simulation. It is the exact segregated pressure-based solver with SIMPLE as the pressure-velocity coupling. A Second Order Upwind scheme was used for continuity, momentum, energy, turbulent dissipation, and kinetic energy.
@@ 4.5.2 Contour of Volume fraction(Liquid)
The blue fill indicates a complete vapour phase, and the red fill indicates a fully liquid phase. The yellow region corresponds to a vapour fraction of 0.6, which indicates a thin two-phase regime, which is a point where the vapor is ready to convert back to the single-phase liquid. It is evident from the above figures that there is a distinct cavitation regime from the throat section up to a certain length in the downstream section of the venturi that corresponds to the Cavitating length of the Venturi.
@@ 5 Experimental Observations
This section puts forth the experimental observations of two experiments, one with the acrylic venturi test section and the other with the aluminum venturi test section. The acrylic venturi was incompatible with the cryogenic temperature flow conditions, resulting in the crack of the test section itself.
@@ 5.2 Observations from experiment - 2
The figure 53 shows a distinct vapor regime, which starts from the throat region and extends to some regions in the diverging section of the venturi. The reading from the pressure measurement indicated a near-cavitation operating condition.
Figure 54: Sample cavitating length measurement
The figure 53 was image processed to get the cavitation length, shown in the figure 54.
The following figure 55 shows the images and corresponding cavitating length for the other test cases.
@@ 6 Conclusion
  1. 1.
    The attempted numerical 1D model based on the extension of Wang and Brennen [9] model with an additional thermal term in line with numerical 1D model of Rodio [8] was implemented using MATLAB code. Following are the conclusions drawn from the current 1D numerical studies:
    • A benchmark study of this model with water was done, and the results matched well with Wang and Brennen [9] model. The superimposed plots of the current model with the actual plot of Wang and Brennen [9] are shown in the figures4,5,6.
    • It was also interpreted that the thermal term in a cavitating water flows had a negligible role to play in bubble growth or collapse as the plots for the magnitude of source terms 10 &11had an order of 10-1 for the thermal term and 104 for pressure term.
    • However, the present model has shortcomings in predicting realistic performance parameters of cavitating venturi.
  2. 2.
    The 2D numerical model was implemented in ANSYSFluent, and the following are the outcomes of the simulation studies.
    • The cavitating length predicted for benchmarking the experimental results of hydrogen cavitation for different test cases of Hord et al. [5] whose details are shown in the figures 37,38 for validating the developed 2D numerical model.
    • A similar procedure was extended for the actual symmetric venturi using liquid nitrogen using the experimental data as the boundary conditions as shown in the figure 43, and the cavitating length and errors for the same with experimental cavitating lengths were predicted as detailed in the figures48 &49, which showed an error of less than $5\%$ that could be accepted as reasonable predictions.
    The following are the outcomes from the Experimental observations
    • The acrylic venturi test section did not withstand the low-temperature liquid nitrogen flow, resulting in leaks near the upstream venturi and altering the flow pattern to stratified conditions as shown in 52. Hence, a distinct cavitating flow could not be observed in experiment 1.
    • The experiment-2, which employed an aluminum venturi, showed distinct cavitating regimes when operated at an initial dewar pressure of 1.5 bar as shown by the figure 53
    • The cavitating length measured from the experimental image is shown in the figures54 & 55
@@
The above algorithm is incorporated in Matlab coding, and the model was solved for cases involving water and liquid nitrogen. Numerical solutions with water after including the thermal model are discussed in the following section.

### v2

@@ 2.5.1 Case-1
The void fraction plot, as shown in figure 14, increases continuously up to 3.8 ×$\times 10^{-4}$ after the nozzle exits from the location of 2.7 to the far downstream of the nozzle to reach the value of 1 which corresponds to fully vapour condition. From the plot for the coefficient of pressure shown in the figure 15, the same behaviour is inferred.
@@ 2.5.2 Case-2
Figure 19: Non-dimensional bubble radius plot for the case-2 of Liquid Nitrogen
@@ 2.5.3 Case-3
The 1D numerical model for cavitating flows with the thermal effects resulted in successful benchmarking of the Wang &\&& Brennen [9] solution for water. The extension of the same model for cryogenic cavitating nozzle flows resulted in the damping solutions as the behaviour of cryogenic liquid nitrogen distinctly differs from the water. Overall, it can be ascertained that the 1D numerical model had limited predictability of the nature of cavitating flows, which are restricted to bubble radius, void fractions, coefficient of pressure, and the thermal and pressure source terms corresponding to a single bubble as the governing Rayleigh Plesset equation itself is valid for dynamics of the single bubble only. However, cavitation is a cluster phenomenon that demands the extension of the studies to two-dimensional numerical and experimental studies, whose attempted results are given in the subsequent sections.
@@ 3 Two-dimensional model CFD Studies
The following section systematically starts with an overview of the methodology adopted for 2D numerical studies using the commercial CFD solver ANSYS Fluent. The simulations have been carried out to capture the cavitating flows in a Venturi profile chosen to be symmetric, with a turbulence model to get insight into the actual experimental work. Simulations were done for cryogenic liquid nitrogen. The governing equations, such as continuity and momentum equations for the multiphase model, are solved using finite volume fluent code. Here, we discusses the viscous modeling, interfacial effects, solution methodology, and the boundary conditions used for the benchmarking test case of Hord et al. [5] and the extension of the same for current 2D numerical work.
@@ 3.1 Turbulence Modelling
The actual cavitating nozzle and venturi flows are turbulent in nature. Hence, a standard turbulence model has been incorporated into the simulations. Realizable $k-\varepsilon$ was used in simulations. The chosen model of $k-\varepsilon$ is in line with the internal flows, and the coefficients for the same are the standard values built into the ANSYS-fluent. The model has two transport equations, one for turbulent kinetic energy k and another for turbulent dissipation $\varepsilon$. The two transport equations are given as follows
@@
The turbulent viscosity $\mu_{t}=\rho C_{\mu}\frac{k^{2}}{\varepsilon}$, where $C_{\mu}$ is a function of mean strain rate, the angular velocity of the system rotation, and the turbulence fields such as k and $\varepsilon$. The model constants are given by,
@@ 3.2.1 Schnerr Sauer model for Mass Transfer
Rb is the bubble radius, and $\eta_{b}$ is the bubble population number per unit volume of the liquid.
This model assumes that the number of bubbles is neither created nor destroyed. To model bubble dynamics, the Rayleigh-Plesset equation in simple form was incorporated that was built with Schnerr-Sauer model, which is given by
@@
Where Pb is the bubble pressure, P is the pressure, far away from the bubble. Using the above relations, the mass source term is
@@ 3.3 Interfacial Heat Transfer Effects
3.4 Benchmark of 2D simulations with experimental work of Hord et al.
An attempt was made to benchmark the 2D numerical work with the experimental test case of Hord 121B [5] for validating the cavitating length for liquid hydrogen flows in a venturi profile.
A finite-volume mesh using Ansys workbench was performed for the actual venturi profile used by Hord. A grid independence study was done for the venturi, and a grid- independent solution was obtained for a mesh with 51,335 nodes and 50527 elements. The following figure shows the meshed model of the venturi profile.
@@
The simulation was done with the boundary conditions taken from the experimental data of Hord et al. [5].
@@ 3.4.1 Solution Methodology
The simulations were performed using fluent with the Eulerian multiphase model. An implicit finite-volume scheme is based on a segregated pressure-based solver with SIMPLE as the pressure-velocity coupling. A second-order upwind scheme was used for continuity, momentum, energy, turbulent dissipation, and kinetic energy.
@@
The residuals were set to 0.00001 for continuity, velocity, K, and $\varepsilon$. For the energy equation and volume fraction, the residuals were set to be 1-06.
@@ 3.4.2 Simulation Results
The following figure32 shows the simulation results of Hord test case 121B. The blue fill indicates the complete liquid phase, and the red fill indicates the complete vapour phase.
@@ 3.4.3 Estimating Cavitating length
Similar work was done by Rodio et al. [8], and the contour of the void fraction of the vapour obtained for the Hord test case 121B [5] is shown in the following figure.
@@
It was observed that the simulation results by Rodio et al. [8] overpredicts the experimental cavitating length observed by Hord et al. [5]. The following measurement proofs validate the above interpretation of the result.
@@
The calculated cavitating length from the above contours of volume fraction gives a length of 11.8235 cm compared to the actual experimental cavitating length of 8.26 cm, measured by Hord et al. [5]. However, the wrong prediction of cavitation length may not be precisely due to the model used, as mesh methods and chosen wall functions have significant influences on the convergence of numerical solution. Having benchmarked the 2D numerical model for cryogenic cavitating venturi with Hord et al. [5] experimental test cases for predicting the cavitating length in the venturi profile. The same is now extended for the current work, the details of which are given in the following section.
@@ 4.1 Computational Model Details
A steady-state, two-dimensional, incompressible, turbulent Eulerian mixture model was developed based on finite-volume formulation. The details of the computational model are shown below
@@
The figure41depicts the actual 2D model of the venturi region of the test section developed for simulations. The dimensions correspond to the actual dimensions of the experimental setup. The boundaries include an inlet, wall, symmetry, and the outlet.
@@ 4.2 Boundary conditions
The outlet boundary condition was atmospheric pressure, as the experiment’s venturi profile opens to the atmosphere. Liquid temperature condition at the outlet is assumed to be 77 K, and vapour fraction at the outlet is assumed to be 0 indicating a single-phase liquid.
@@ 4.3 Solution Methodology
The same solution methodology was used for the Hord benchmark simulation. It is the exact segregated pressure-based solver with SIMPLE as the pressure-velocity coupling. A second-order upwind scheme was used for continuity, momentum, energy, turbulent dissipation, and kinetic energy.
@@ 4.5.2 Contour of Volume fraction(Liquid)
The blue fill indicates a complete vapour phase, and the red fill indicates a fully liquid phase. The yellow region corresponds to a vapour fraction of 0.6, which indicates a thin two-phase regime, which is a point where the vapour is ready to convert back to the single-phase liquid. It is evident from the above figures that there is a distinct cavitation regime from the throat section up to a certain length in the downstream section of the venturi that corresponds to the cavitating length of the venturi.
@@ 5 Experimental Observations
This section puts forth the experimental observations of two experiments, one with the acrylic venturi test section and the other with the aluminium venturi test section. The acrylic venturi was incompatible with the cryogenic temperature flow conditions, resulting in the crack of the test section itself.
@@ 5.2 Observations from experiment - 2
The figure 53 shows a distinct vapour regime, which starts from the throat region and extends to some regions in the diverging section of the venturi. The reading from the pressure measurement indicated a near-cavitation operating condition.
The following figure 54 shows the images and corresponding cavitating length for the test cases.
@@ 6 Conclusion
The attempted numerical 1D model based on the extension of the Wang and Brennen [9] model with an additional thermal term in line with the numerical 1D model of Rodio [8] was implemented using a finite-difference code in MATLAB. It was inferred that the thermal term in a cavitating water flow had a negligible role to play in bubble growth or collapse. On the other hand, this thermal term had a significant influence on the bubble radius for the liquid nitrogen case. The 2D numerical model was implemented in ANSYS Fluent, and the following are the outcomes of the simulation studies.
• The cavitating length predicted and compared with the experimental results of Hord et al. [5] for validating the developed 2D numerical model.
• A similar procedure was extended for the chosen cavitating venturi using the experimental data as the boundary conditions, and the cavitating length was predicted for the same.
The following are the outcomes from the experimental observations:
• The acrylic venturi test section did not withstand the low-temperature liquid nitrogen flow, resulting in leaks near the upstream venturi and altering the flow pattern to stratified conditions. Hence, a distinct cavitating flow could not be observed in experiment 1.
• The experiment-2, with an aluminum venturi, showed distinct cavitating regimes and the cavitating lengths were estimated from the recorded high-speed camera experimental images.
@@
The above algorithm is incorporated in Matlab coding, and the model was solved for cases involving water and liquid nitrogen.